OpenStudy (anonymous):

NEED HELP SOLVING THIS ! :C Karl owes $7,464 on a credit card with a 20.4% interest rate compounded monthly. What is the monthly payment he should make in order to pay off this debt in 12 months, assuming he does not charge any more purchases with the card?

OpenStudy (tkhunny):

i = 0.204 j = i/12 v = 1/(1+j) 7464 = Pmt(v + v^2 + v^3 + ... + v^12) = \(Pmt*\dfrac{v-v^{13}}{1-v}\) It's a present value problem. They all work the same when contemplated through Basic Principles.